Let f(n)=2cosnx​​ Anbelons to N then f(1)f(n+1)-f(n) is equal to

  1. Let f(n)=2cosnx​​Anbelons to N then f(1)f(n+1)-f(n) is equal to

Grade:11

1 Answers

Arun
5 years ago
f(1) = 2 cosx
f(n+1) = 2 cos(n+1)x
f(n)  = 2 cos nx
 
now 
 
f(1) f(n+1) = 2 [cos (n+2)x + cos nx]
 
hence
 
f(1) f(n+1) – f(n) = 2 cos (n+2)x
 
Hope it helps
 
Regards
Arun (askIITians forum expert)